In a gas turbine engine, the compressor takes in air at a temperature of 15°C, pressure of 100 kPa, and a volumetric flow rat(f) Find the net power generated by the engine.
Answer #1

Drawing the.T-s diagram and then finding Every point temperature we can solve this question as below-Sol - The data provided in the question are as below - Inlet Pressure to compressure -100 kpa 15c tempo = 289 k. PA T 5 m3/sProcess 1-2:- compression . A-3:- constant Pressure heat addition. 3-41:- Expansion . bol:- constant pressure heat rejectionIs entropic efficiency of compressor, Qis) C Ideal work! Actual work = Tz - Ti torti 406-233 Ta 288 Oy, 0.82= -288 or, Tz0.78 - = or, Ta 432186 695.140 – 432.186 oy, (TQ 637 290 K Ans (6) There fore , temperature at the exhaust exhaust of the hea56.3594 x lil x (432 • 186 –288) 1008.63 KW. Turbine work, WT = mcp (T3 - Tu) = 6:3594 x 1+1 X (873-6950140) 1244.991 KW, =f Net power generated by the engine, whet =Wy -We (12.441191 - 1008563 Il ) KW ㄹ 235.561 kw Ans. 6).